A 180-day note dated on March 14, 2019 is made at 9.25% for $1665.00. what is
the present value of the note on May 26, 2019 if the money is worth 10.5%?
Answer #1

Issue price of 9.25% worth of note = $1665

Redemption price of note = 1665 + (1665*9.25%*180/360) = 1665 + 77 = $ 1,742

On may 26, 2019, after 72 days from Mar 14, 2019, Balance no. of days left = 180-72 = 108 days

Interest rate = 10.5%

PV of Note = 1742 / (1+10.5%*108/360) = 1742 / (1+0.0315) = 1742 / 1.0315 = $ 1,688.80

The present value of note on may 26, 2019 is $ 1,688.80.